### v1

@@ 48 Work of partitioning, Hill free energy, and subdivision potential
We consider a system initially in a stable equilibrium state and examine the thermodynamic cost of partitioning it into $\lambda$ compartments. Each compartment is required to contain equal amounts of constituents, equal volume, and equal entropy, so that the partitioned system as a whole is again in a stable equilibrium state. Since the overall energy and entropy of the system are conserved in a reversible weight process, the transition from the unpartitioned to the partitioned configuration can, in principle, be carried out reversibly, and therefore admits a well-defined minimum work requirement.
@@
It is worth emphasizing that the existence of a nonzero minimum work of partitioning reflects a departure from extensivity. In macroscopic thermodynamics extensivity is assumed as part of the simple-system model approximation Gyftopoulos and Beretta (1991); Beretta and Gyftopoulos (2015)), and it implies that subdivision into identical parts carries no energetic cost. For systems with few particles, this approximation fails: wall rarefaction and surface effects, finite-size constraints, and changes in the spectrum of accessible states introduce additional energetic contributions associated with subdivision. The Hill (Euler) free energy quantifies precisely this contribution and is therefore also referred to as the subdivision potential. Its appearance here highlights the fact that thermodynamic descriptions applicable to systems of arbitrary size must explicitly account for deviations from extensivity, rather than assuming them away from the outset. This is made even more explicit in the next section.
@@ 51 Entropy and uncertainty in quantum models
There are numerous interesting results that stem from the application of the concepts discussed so far to the study of additional properties defined within the framework of quantum-theoretical models. This topic serves as an example illustrating how the foregoing exposition of thermodynamic concepts connects with the foundations of quantum physics, which over the past few decades have become the basis of a wide range of modern technologies. It also allows us to explore aspects such as the interpretation of entropy as a measure of uncertainty and the origin of the ideal-gas model.
@@
In addition to quantization, the aspect that most sharply contrasts with the determinism on which classical physics was based is the unpredictability of the outcome of a single measurement act. This unpredictability reflects an intrinsic indeterminacy in the state of systems and is a defining discovery of modern physics, its technological applications, and the various philosophical interpretations that have emerged over the last century and remain under discussion.606060“God does not play dice!” declared Albert Einstein—one of the founders of quantum theory—expressing his profound dissatisfaction with the impossibility of predicting which energy level a system will yield in a given measurement act. Today, this feature is widely accepted as a fact of physical reality, although the dissatisfaction expressed by Einstein is still shared by some.
@@
where $k$ is the Boltzmann constant, ${k_{\text{\scriptsize B}}}=1.38066\times 10^{-23}$ J/K, and $p_{\epsilon_{j}}$} and $g_{\epsilon_{j}}$} are the probabilities and multiplicities of all energy levels.

### v2

@@ 48 Work of partitioning, Hill free energy, and subdivision potential
The partitioning of a system into subsystems is a fundamental operation that clarifies the distinction between additivity and extensivity. Macroscopic and mesoscopic treatments often take for granted the “simple-system model approximation” Gyftopoulos and Beretta (1991); Beretta and Gyftopoulos (2015), whereby a system in a stable equilibrium state can be subdivided into a composite of smaller subsystems without energetic or entropic cost. This approximation is a cornerstone of the continuum hypothesis required to define spatial fields in many nonequilibrium frameworks; it amounts to neglecting wall rarefaction, surface effects, finite-size constraints, and changes in the spectrum of accessible states introduced by physical partitions.
For systems with few particles, however, this approximation fails. Such effects introduce non-negligible energetic and entropic contributions associated with the subdivision of a system into subsystems. Here, it is vital to emphasize that “subsystems” refer to well-defined systems (separable and uncorrelated) — a fundamental condition for their individual energies and entropies to be formally defined.
As part of our rigorous operational construction, which remains valid for both few- and many-particle systems, we must explicitly account for the work required to establish the boundaries that define a partition. As we shall see, the Hill (Euler) free energy quantifies precisely this contribution and is therefore referred to as the subdivision potential. Its appearance in our derivation highlights that a thermodynamic description applicable to systems of arbitrary scale must account for deviations from extensivity rather than assuming them away from the outset — a necessity made even more explicit in the following section.
We consider a system initially in a stable equilibrium state and examine the thermodynamic cost of partitioning it into $\lambda$ compartments. Each compartment is required to contain equal amounts of cconstituents, equal volume, and equal entropy, so that the partitioned system as a whole is again in a stable equilibrium state. Since the overall energy and entropy of the system are conserved in a reversible weight process, the transition from the unpartitioned to the partitioned configuration can, in principle, be carried out reversibly, and therefore admits a well-defined minimum work requirement.
@@ 51 Entropy and uncertainty in quantum models
The following discussion of Quantum Thermodynamics serves as a concrete demonstration of the operational construction developed in the first part of this work for one of the nonequilibrium frameworks listed in Section 5. In the quantum regime, the identification of the thermodynamic state and its properties is particularly challenging, as many classical heuristic partitions of energy transfer into work and heat are currently the subject of active investigation Binder et al. (2018); Davoudi et al. (2024). There are numerous interesting results that stem from the application of the concepts discussed so far to the study of additional properties defined within the framework of quantum-theoretical models. Our framework provides a self-consistent logical scaffolding that remains independent of specific modeling choices for dissipative processes and may help to clarify currently controversial definitions. This discussion also allows us to explore fundamental aspects such as the interpretation of entropy as a measure of uncertainty and the origin of the ideal-gas model.
@@
In addition to quantization, the aspect that most sharply contrasts with the determinism on which classical physics was based is the unpredictability of the outcome of a single measurement act. This unpredictability reflects an intrinsic indeterminacy in the state of systems and is a defining discovery of modern physics, its technological applications, and the various philosophical interpretations that have emerged over the last century and remain under discussion.616161“God does not play dice!” declared Albert Einstein — one of the founders of quantum theory — expressing his profound dissatisfaction with the impossibility of predicting which energy level a system will yield in a given measurement act. Today, this feature is widely accepted as a fact of physical reality, although the dissatisfaction expressed by Einstein is still shared by some.
@@
where $k$ is the Boltzmann constant, ${k_{\text{\scriptsize B}}}=1.380649\times 10^{-23}$ J/K, and $p_{\epsilon_{j}}$} and $g_{\epsilon_{j}}$} are the probabilities and multiplicities of all energy levels.